OpenStudy (anonymous):

simplify. 7x+3y-2+6x-1+y^2

alones (alones):

Hey there Welcome to Openstudy!!

OpenStudy (anonymous):

hi

alones (alones):

Let's simplify step-by-step this equation 7x+3y−2+6x−1+y2 =7x+3y+−2+6x+−1+y2 don't forget to combine Like Terms =7x+3y+−2+6x+−1+y2 =(y2)+(7x+6x)+(3y)+(−2+−1) =y2+13x+3y+−3 Hope that helps ! :)

OpenStudy (anonymous):

thank you!!!

alones (alones):

You welcome<3 :)

alones (alones):

btw don't forget to close oyur question too :) Best Wishes